Write the Integrating factor of the Differential Equation dy/dx
+ x/1+x *
y=ex
1
Expert's answer
2020-06-16T15:05:53-0400

"\frac {dy}{dx} + \frac {x}{1+x} y = e^x"

Comparing with "\frac {dy}{dx} + Py = Q" we get

P = "\frac {x}{1+x}" = "\frac {x}{x+1}" and Q = ex

So integrating factor is

"e^{\int Pdx}" = "e^{\int \frac {x}{x+1}dx}" = "e^{\int \frac {x+1-1}{x+1}dx}"

= "e^{\int ( \frac {x+1}{x+1} - \frac {1}{x+1})dx}"

= "e^{\int ( 1 - \frac {1}{x+1})dx}"

= "e^{x - ln(1+x)}"

= "\frac {e^x}{e^{ln(1+x)}}"

= "\frac {e^x}{1+x}"

"\mathbf {Answer}"

The integrating factor is "\frac {e^x}{1+x}"